2. The weight of potato chips in a medium-size bag is stated to be 10 ounces. The amount that the packaging machine puts in these bags is believed to have a Normal model with mean 10.2 ounces and standard deviation 0.12 ounces. Some of the chips are sold in “bargain packs” of 3 bags. What’s the probability that the mean weight of the 3 bags is below the stated amount?

Please show all work.

Please no handwritten work, just typed.
Answer #1

GIVEN:

mean=10.2

standard deviation=0.12

n=3

the probability that the mean weight of the 3 bags below the stated amount

p(xbar less than 10)= P(X < 10)

=p(x-µ/σ/N < 10-10.2/0.12/ √ 3)

= p(z<-0.2/0.07)

=p(z<-2.85)

=1-p(z<2.85)

= 1-0.0978

=0.0022